Rate constant of first order reaction A- products is 0.016/min calculate time required for 80%reaction to be completed

Grade:12th pass

1 Answers

Vikas TU
6 years ago
For the principal arrange response: 
k = (2.303/t) log (a/a – x) t= ?, when x = 0.8 a 
From above, k = 0.016 min-1 
Along these lines, t = (2.303/0.016) log (a/a – 0.8 a) 
= (2.303/0.016) log (1/0.2) = 1.58437 min 
The time at which the response will be 80% finished is 1.58437 min.